MATHS:ellipse Find the equation of the ellipse with the centre at the origin and whose focii are (0,±3) & e=3/4.

MATHS:ellipse
Find the equation of the ellipse with the centre at the origin and whose focii are (0,±3) & e=3/4.

Grade:11

1 Answers

Vikas TU
7 years ago
Focus point is given as: = > ae = 3
3a/4 = 3 = > a = 4
and
from
e = root(1 – (b/a)^2) = > ¾
1 – 9/16 = (b/a)^2
b/a = root(7)/4
b = 3*root7/4
Eqn. of ellipse = >
x^2/16 + y^2/(63/4) = 1